OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-2, 2) and is parallel to y = 4x + 7.

OpenStudy (anonymous):

A: y=4x+10 B: y=1/4x + 5/2 C: y=-4x-6 D: y= -1/4x + 3/2

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

y-2/x+2 = 4 [Coz point (-2,2) is given and slope = 4]

OpenStudy (anonymous):

Simplify and get ur result

OpenStudy (anonymous):

i dont remember how to do any of this

OpenStudy (anonymous):

I DONT KNOWWWWW PLEASE WALK ME THROUGH THIS PLEEEAASEE

OpenStudy (anonymous):

i learned this so long ago

OpenStudy (anonymous):

y - 2 = 4x + 8 so, y = 4x + 10

OpenStudy (anonymous):

OHHH

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

YW